OpenStudy (anonymous):

11 minus the quotient of a number and 8 is 15? I need help in writing it

OpenStudy (anonymous):

11-(-4)=15

OpenStudy (anonymous):

11-x/8=15

OpenStudy (anonymous):

from: 11-x/8=15

OpenStudy (triciaal):

quotient of a number and 8 let the number = x quotient what you get when you divide x/8 11 minus (quotient of a number and 8) = 15 minus means subtract is means equal 11-x/8 = 15